OpenStudy (anonymous):

Solved the system by substitution.if the system is inconsistent and has no solution,state this .if the system is dependent,write the form of the solution for any real number x. 9x-6y=19 -x+8y=4

OpenStudy (anonymous):

Shakur, is this the same question?

OpenStudy (anonymous):

X=8y-4 9(8y-4)-6y=19 72y-36-6y=19 66y=55 Y=5/6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ DSS so your answer a solution or has no solution or inconsistent or depedemt